OpenStudy (anonymous):

Angle DEF and Angle FEG are complementary. m angle DEF=(3x-4) and m angle FEG= (5X+6)

OpenStudy (dayakar):

angleDEF+angleFEG= 90 (complementary angles)

OpenStudy (dayakar):

3x-4+5x+6=90

OpenStudy (dayakar):

8x+2=90

OpenStudy (dayakar):

8x=90-2 8x=88

OpenStudy (dayakar):

x= 88/8 x=11

OpenStudy (dayakar):

if sum of the two angles is equal to 90 degrees then the two angles are called as complementary angles